You need to show the work in order to receive full credit

Formula:

BE point = Fixed Cost divided by (price minus unit variable cost)

Given

Quantity Sold = 250 units

Fixed Cost = $1,500

Unit variable cost = $9 per unit

Unit Price = $12

1) Calculate for break event point

2) If the company making profit or loss?

3) How much profit or loss?
Answer #1

1) Break-Even Point :- Fixed Cost / (Price -Unit Variable cost)

BE = 1500/(12-9) = 1500/ 3 = 500 units

2) Company is making loss ,since they have sold below break even.

3) Total Sales = Quantity Sold * Price = 250 * 12 = $3000

Total Profit = Sales - Variable Cost -Fixed Cost

Total profit = 3000 - 2250 - 1500 =$ (-750)

Alternatives ,

Loss = ( Break even units - Units Sold ) * Contribution = (500-250)*3 = $750